Geometry. The length of a rectangle is 2 inches more than twice its width. If the perimeter of the rectangle is 34 inches, find the dimensions of the rectangle.



Answer by malakumar_kos@yahoo.com(315)


The length of a rectangle is 2 inches more than twice its width. If the perimeter of the rectangle is 34 inches, find the dimensions of the rectangle
Let the width of the rectangle = w = x inch
the length of the rectangle = l = 2x+2 inch
perimeter of the rectangle = p = 34 inch
p = 2(l+w)
34 = 2(2x+2+x)
34/2 = 3x+2
17-2 = 3x
15 = 3x
x = 15/3 = 5
width = 5 inch
length = 2*5+2 = 10+2 = 12 inch